OpenStudy (anonymous):

first correct answer will get a fan and a medal!!!! solve, using the substation method. a.) there are an infinite number of solutions. b.) (12,30) c.)(14,12) d.) there is no solution

OpenStudy (anonymous):

what is the question

OpenStudy (anonymous):

oops its y=2x+6 3x-y=6

OpenStudy (anonymous):

ok so 3x-(2x+6)=6

OpenStudy (anonymous):

if you solve that for x, you'll find both x and y, and your answer.

OpenStudy (anonymous):

\[\large \color{red}{y}=\color {green}{2x+6}\] \[\large 3x-\color {red}{y}=6~becomes~ 3x-(\color {green}{2x+6}) =6\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

do you need help with solving for x?

OpenStudy (anonymous):

.umm nope I got it thank you so much

OpenStudy (anonymous):

alright